1- Suppose a firm issues a single bond with a face value of $100,000 which is Convertible into 10,000 Ordinary Shares with a Par Value of $1. The credit entry to “Share Premium – Conversion Equity” on the date of issue was for $20,000.   Assume that the maturity date of the bonds has now arrived and the bonds will be converted into Ordinary Shares.

The Debit Entry to the “Bonds Payable” account on the date of conversion will be for an amount of:

2-: Suppose a firm earns Net Income of $1,200,000. The company pays an Ordinary Dividend of $400,000 and a Preference Dividend of $200,000. Throughout the financial year, the firm has 500,000 Ordinary Shares and 200,000 Preference Shares. The firm’s Earnings Per Share (EPS) is:

3-: Suppose a firm earns Net Income of $1,000,000. The company does not pay dividends.

At the start of the financial year the firm had 530,000 Ordinary Shares. On 31 March, the firm issued 80,000 Ordinary Shares for Cash. On 30 November, the firm issued a further 120,000 Ordinary Shares for Cash.    

The number of Ordinary shares used in the calculation of Earnings Per Share (EPS) is:
Answer #1

Part 1)

The correct answer bond payable will be debited With $ 100000

Explanation

The difference between bond value and market price of the shares (100000-(3*10000)) will be recognized as income and bond payable will be credited with $ 100000

part 2)

The correct answer is Earning per share = $ 2 per share

EPS = (net income - preference dividend) / common shares

= (1200000-200000)/500000

=$ 2 Per share

Part 3)

The correct answer is number of ordinary shares for Earning per shares 600000 Shares

Weighted average shares = beginning shares + (shares issued on March 31 *9/12) + (Shares issued on November 30 *1 /12)

= 530000+(80000*9/12)+(120000*1/12)

= 530000+ 60000 +10000

= 600000 Shares